how to find a line that passes through the points (-10, -4) and (-1,2)

Answer by Alan3354(69443)

how to find a line that passes through the points (-10, -4) and (-1,2)
===========
You can use determinants:
---
|+x +y + 1|
|-10 -4 1| = 0
|-1 +2 +1|
---
x*(-4-2) - y*(-10+1) + 1*(-20-4) = 0
-6x + 9y = 24
2x - 3y = -8
-------